There were 82 people at a family reunion.
There were 4 more men than women.
There were 10 more children than adults.
How many childern were there?

Answer by edjones(7311)

w=number of women, w+4=numb of men, w+w+4+10=numb of kids
w+w+4+w+w+4+10=82
4w+18=82
4w=64
w=16
w+w+4+10=46
.
Ed